OpenStudy (anonymous):

What is the distance between points P(10, 7) and Q(1, 8)?

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) does this help ??

OpenStudy (anonymous):

yah it wold be square root of 73

hartnn (hartnn):

73 ?? how ?

OpenStudy (anonymous):

wait sorry that wrong my bad

hartnn (hartnn):

(x1-x2)^2 = (10-1)^2 =....?

OpenStudy (anonymous):

thats 81

hartnn (hartnn):

ok, so tell me what you get after redoing it....

hartnn (hartnn):

right, and (y1-y2)^2 = (7-8)^2 =... ?

OpenStudy (anonymous):

1

hartnn (hartnn):

yup! \(\sqrt{81+1}=..?\)

OpenStudy (anonymous):

81

hartnn (hartnn):

81+1=82 so, \(d=\sqrt {82}\)